A gaseous mixture consists of 1.27 mol of helium, 0.36 mol of neon, and 1.58 mol of argon.
What is the mole fraction of helium in the mixture?
1
Expert's answer
2019-12-09T08:04:40-0500

W(He) = 1,27/(1,27+0,36+1,58) = 1,27/3,21 = 0,3956 or ≈ 40%